Suppose that 40% of a large population are in favor of candidate A. A pollster selects a random sample of 20 people and asks them about their opinion about the candidate. Let X be the number of people in favor of candidate A. What is E(X) and Var(X)?
Answer #1

This is a binomial distribution.

n = 20

p = 0.4

E(X) = n * p = 20 * 0.4 = 8

Var(X) = n * p * (1 - p) = 20 * 0.4 * (1 - 0.4) = 4.8